OpenStudy (anonymous):

What is the slope of the line that passes through the points (–9, 5) and (–7, 10)?

pooja195 (pooja195):

y2-y1 x2-x1

OpenStudy (anonymous):

use the formula

pooja195 (pooja195):

Plug the numbers in

OpenStudy (anonymous):

\[slope~m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

can u plug in the values? @zkmath

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

10-(-5)

OpenStudy (anonymous):

5 @MARC_

pooja195 (pooja195):

thats not right

OpenStudy (anonymous):

\[=10-(-5)\]\[=10+5\]\[=?\]

OpenStudy (anonymous):

\[slope~m=\frac{ 15 }{ -7-(-9) }\]

OpenStudy (anonymous):

\[slope~m=\frac{ 15 }{ 2 }\]

OpenStudy (anonymous):

simplify i dont know how

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

u cant simplyfy that =/

OpenStudy (anonymous):

well no

OpenStudy (anonymous):

use ur calculator

OpenStudy (anonymous):

what do you tupe in

OpenStudy (anonymous):

15divide2

pooja195 (pooja195):

you would get a decimal

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

7.5 thats easy

pooja195 (pooja195):

right

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Therefore,the slope of the line that passes through the points (–9, 5) and (–7, 10) is 7.5

OpenStudy (anonymous):

but there is no answer like tht

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

are there any choices? @zkmath

OpenStudy (anonymous):

yes there are

OpenStudy (anonymous):